# Taming the Tri-Space Tension: ARC-Guided Hallucination Modeling and Control for Text-to-Image Generation

### v1

@@ Abstract
Despite remarkable progress in image quality and prompt fidelity, text-to-image (T2I) diffusion models continue to exhibit persistent ”hallucinations”, where generated content subtly or significantly diverges from the intended prompt semantics. While often regarded as unpredictable artifacts, we argue that these failures reflect deeper, structured misalignments within the generative process. In this work, we propose a cognitively inspired perspective that reinterprets hallucinations as trajectory drift within a latent alignment space. Empirical observations reveal that generation unfolds within a multiaxial cognitive tension field, where the model must continuously negotiate competing demands across three key critical axes: semantic coherence, structural alignment, and knowledge grounding. We then formalize this three-axis space as the Hallucination Tri-Space and introduce the Alignment Risk Code (ARC): a dynamic vector representation that quantifies real-time alignment tension during generation. The magnitude of ARC captures overall misalignment, its direction identifies the dominant failure axis, and its imbalance reflects tension asymmetry. Based on this formulation, we develop the TensionModulator (TM-ARC): a lightweight controller that operates entirely in latent space. TM-ARC monitors ARC signals and applies targeted, axis-specific interventions during the sampling process. Extensive experiments on standard T2I benchmarks demonstrate that our approach significantly reduces hallucination without compromising image quality or diversity. This framework offers a unified and interpretable approach for understanding and mitigating generative failures in diffusion-based T2I systems.
@@ 1 Introduction
Figure 1: Visualizing hallucination as trajectory drift in latent alignment space. Successful generations (middle row) follow a coherent sampling trajectory from noise to data, remaining close to the prompt intent. Hallucinatory generations (bottom row) exhibit tension drift, leading to semantic and structural deviations.
@@
We begin by revisiting a foundational question in the generative process of diffusion models: Are hallucinations merely stochastic artifacts of sampling, or do they reflect deeper structural failures rooted in internal tension dynamics? To eliminate ambiguity and isolate the source of failure, we construct a synthetic dataset in which each text prompt deterministically maps to a unique image (see Sec. 1.1). Surprisingly, even under these nearly ideal conditions, diffusion models consistently produce significant misalignments, ranging from shape swaps and color inversions to positional inconsistencies. Crucially, these failures are not random. Unsupervised clustering of hallucinated outputs reveals a remarkably stable structure: across diverse prompts and generations, errors consistently segregate into three well-separated categories:semantic drift, structural distortion, and factual misrepresentation. What initially appears as scattered output anomalies now unveils an intrinsic structure of failure modes, hinting at latent alignment axes embedded in the generation process.
A closer examination reveals that these hallucination types correspond to three fundamental cognitive trade-offs the model must continually negotiate: (1) Semantic Coherence: maintaining identity fidelity (e.g., generating a “puppy” rather than a “cat”); (2) Structural Alignment: preserving spatial and compositional correctness (e.g., avoiding positional swaps or shape mismatches); (3) Knowledge Grounding: ensuring factual and commonsense plausibility (e.g., distinguishing a “blanket” from a “carpet”). These findings challenge the traditional view of hallucinations as local attention failures or sampling noise. Instead, we argue they are symptomatic of a deeper phenomenon: the emergence of directional cognitive tension within the generative trajectory.
@@
From this insight, we arrive at a pivotal conceptual shift: Text-to-image generation is not a static mapping but a dynamic traversal through a latent tension space. Each prompt implicitly defines a multiaxial cognitive tension field, and the diffusion process must iteratively traverse this space while balancing competing alignment forces. This dynamic process, termed cognitive alignment tension, unfolds throughout the denoising trajectory. When one or more tension axes dominate, they destabilize the generative equilibrium, causing the trajectory to drift off course. We denote this directional deviation as a trajectory drift $\Delta\vec{t}$, which manifests as a hallucination in the output. Ideally, the generation should evolve within a well-balanced region of latent space, or the ideal alignment manifold $\mathcal{M}_{\text{ideal}}$. However, the imbalance in alignment tensions disrupt this path, pulling the model away from intended semantics.
This realization leads us to a formal abstraction. Despite the seemingly diverse nature of hallucinations, their root causes consistently converge along three principal axes of alignment imbalance. We thus introduce the: Hallucination Tri-Space $\mathcal{T}^{3}$, a structured latent space spanned by three orthogonal tension axes:(1) Semantic Coherence (SC): the alignment between prompt entities and generated object categories; (2) Structural Alignment (SA): the fidelity of spatial layout and positional relationships; (3) Knowledge Grounding (KG): the factual and commonsense plausibility of generated content. Rather than merely classifying hallucinations, $\mathcal{T}^{3}$ enables explicit modeling of tension dynamics. By tracking the directional shift $\Delta\vec{t}$ within this space, we can quantitatively monitor tension imbalance during generation, offering a principled basis for real-time hallucination diagnosis and control.
@@
To quantify the cognitive alignment tension, we project the trajectory drift $\Delta\vec{t}$ onto the three axes of the Hallucination Tri-Space, yielding a real-time vector we term the Alignment Risk Code (ARC): $\vec{\tau}(p,t)=[\tau_{\text{SC}}(p,t),\tau_{\text{SA}}(p,t),\tau_{\text{KG}}(p,t)]^{\top}$. This ARC vector models the live state of alignment tensions during sampling. Its magnitude reflects the overall cognitive stress, its imbalance indicates the degree of tension asymmetry, and its direction characterizes the dominant source of misalignment. Rather than diagnosing hallucinations after generation, ARC enables in-process monitoring of when and how generative deviation begins to emerge. To leverage this signal for intervention, we introduce the TensionModulator (TM-ARC), a lightweight controller that operates fully within the latent space. TM-ARC interprets ARC dynamics and injects targeted, axis-specific corrections through three specialized submodules: (1) SC-Gate: mitigates semantic drift; (2) SA-Tuner: restores spatial integrity; (3) KG-Augment: reinforces factual grounding. Each submodule is activated with adaptive weights derived from the ARC vector, enabling fine-grained, real-time adjustment of the generative trajectory. This ensures the model remains within the ideal alignment manifold $\mathcal{M}_{\text{ideal}}$ throughout the denoising process, significantly reducing hallucinations without compromising image diversity or quality.
We now look back at the three questions raised at the beginning: Why do diffusion models sometimes turn ”puppies” into ”cats”? Why do such phenomena persist even for simple prompts? Can we understand, model, and control them? Our answer is: hallucinations emerge when the generative trajectory is pulled off-course due to accumulating misalignments in semantic, structural, or factual dimensions during the sampling process. Specifically, ”puppies” may become ”cats” because of their close semantic similarity, and the model’s failure to anchor fine-grained meanings. ”Blankets” may be replaced by ”carpets” when commonsense priors are weak and the model defaults to visually plausible substitutes. Even with simple prompts, these errors persist because the diffusion process is inherently a dynamic tension-balancing trajectory rather than a static semantic mapping. The persistence of hallucinations stems not from prompt complexity but from the accumulation of misalignments across multiple cognitive axes. We model this phenomenon using the ARC vector, which captures rising tensions in dimensions such as knowledge grounding and semantic coherence. Leveraging this signal, our TM-ARC controller actively intervenes during generation, steering the trajectory back toward the ideal manifold $\mathcal{M}_{\text{ideal}}$. Our contributions are threefold:
@@
1.1 Empirical Characterization of Hallucination Patterns in Alignment Tension Space
Before formalizing our framework in the following sections, we first present an empirical study to examine whether hallucinations in T2I models exhibit structured internal patterns. To isolate the generative mechanism from data or language ambiguity, we construct a controlled synthetic dataset with fully deterministic prompt-image mappings.
Synthetic Setup.
We construct a synthetic dataset of 1000 prompt-image pairs, where each prompt unambiguously describes two geometric objects with specific shapes, colors, and spatial positions (e.g., ”A red circle on the left, and a green triangle on the right.”). We train a standard unconditional Denoising Diffusion Probabilistic Model (DDPM) on this dataset using fully unsupervised objectives, removing any external semantic signal or supervision.
Observed Imbalance Patterns.
Despite the simplicity of the dataset, 27.8% of generated samples exhibit misalignments that violate prompt specifications. These include shape swaps, incorrect colors, and spatial reversals. Crucially, each error type can be mapped to one or more cognitive alignment axes (e.g., color errors → SC, position flips → SA, shape errors → KG).
Table 1: Clustering quality comparison across different feature spaces. Silhouette (higher is better) measures intra-cluster compactness vs inter-cluster separation; Calinski-Harabasz (higher is better) evaluates cluster dispersion; Davies-Bouldin (lower is better) penalizes overlapping clusters. Our alignment feature representation yields consistently superior clustering quality across all metrics.
Method Silhouette $\uparrow$ Calinski-Harabasz $\uparrow$ Davies-Bouldin $\downarrow$
PCA 0.41 752.3 1.37
Random 3D Mapping 0.21 389.1 2.10
Alignment Features (Ours) 0.63 1480.5 0.84
Alignment Tension Space Characterization.
To further analyze this structured misalignment, we extract a 3-dimensional alignment deviation vector for each generated sample, reflecting its SC, SA, and KG drift magnitudes. Using unsupervised clustering on these vectors, we find that hallucinated samples form three well-separated clusters. As shown in Table 1, our alignment-based clustering significantly outperforms PCA and random projection baselines across multiple metrics. Figure 2 further visualizes this structured clustering in the hallucination tri-space $\mathcal{T}^{3}$.
2 Related Work
Text-to-Image Diffusion Models
Recent advances in text-to-image (T2I) diffusion models have significantly improved image synthesis quality and compositional controllability. Foundational models such as Stable Diffusion and its variants rely on iterative denoising guided by conditioning inputs, yet faithfully aligning multiple modalities remains a central challenge Wang et al. (2023); Wu et al. (2024). Existing models often struggle when handling complex prompts involving object relations, attributes, or scene layouts. Recent innovations such as SoftREPA Lee et al. (2025), VLAD Johnson & Wilson (2025), and ELBO-T2IAlign Zhou et al. (2025) propose more explicit alignment mechanisms. These include contrastive loss for semantic preservation, hierarchical prompt parsing, and ELBO-based calibration to address subtle alignment errors.
Text-to-Image Diffusion Models
Recent advances in text-to-image (T2I) diffusion models have significantly improved image synthesis quality and compositional controllability. Foundational models such as Stable Diffusion and its variants rely on iterative denoising guided by conditioning inputs, yet faithfully aligning multiple modalities remains a central challenge Wang et al. (2023); Wu et al. (2024). Existing models often struggle when handling complex prompts involving object relations, attributes, or scene layouts. Recent innovations such as SoftREPA Lee et al. (2025), VLAD Johnson & Wilson (2025), and ELBO-T2IAlign Zhou et al. (2025) propose more explicit alignment mechanisms. These include contrastive loss for semantic preservation, hierarchical prompt parsing, and ELBO-based calibration to address subtle alignment errors.
Alignment Control and Conditioning Misalignment
Recent methods attempt to improve condition adherence through attention modulation or conditioning decoupling. For example, Text-Anchored Score Composition (TASC) Wang et al. (2023) explicitly decomposes conditional scores to mitigate attention conflicts, while attention modulation frameworks dynamically adjust cross-attention maps to strengthen alignment fidelity Wu et al. (2024). Although these approaches improve adherence to individual conditions, they do not model holistic hallucination risks across semantic, structural, and knowledge dimensions. To tackle this, alignment-centric frameworks such as CoMat Jiang et al. (2024) and MCCD Li et al. (2025) integrate cross-modal consistency and collaborative scene parsing directly into the generation process. These approaches proactively model alignment gaps, offering more robust compositional generalization during generation rather than relying on post-hoc correction Lu et al. (2025); Lv et al. (2025).
Hallucination Mitigation in Multimodal Generative Models
Hallucination has also been widely studied in large language models (LLMs) and multimodal LLMs. Recent efforts address hallucination detection Chen et al. (2024) and propose retrieval augmentation or consistency supervision for factual grounding Lv et al. (2025); Lu et al. (2025). However, most existing methods operate post-hoc, lacking proactive hallucination prevention during generation. In contrast, our work directly integrates alignment risk modeling into the T2I generation process.
3 Definitions and Preliminaries
We briefly review the T2I diffusion framework that forms the basis of our alignment tension modeling.
Diffusion Process
Given a clean image $x_{0}\sim q(x)$ from the data distribution, the forward diffusion process adds Gaussian noise over $T$ timesteps:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ (1)
where $\beta_{t}\in(0,1)$ controls the noise level at step $t$. This process allows direct sampling from $x_{t}$ via:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (2)
with $\bar{\alpha}_{t}=\prod_{j=1}^{t}(1-\beta_{j})$.
The reverse process learns to denoise through a parameterized model $p_{\theta}$:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ (3)
where $\mu_{\theta}$ is typically computed from predicted noise $\epsilon_{\theta}(x_{t},t)$.
In T2I diffusion, generation is conditioned on text embeddings $c$, modifying the denoising objective to predict $p_{\theta}(x_{t-1}\mid x_{t},c)$.
@@ 4 Why Do T2I Diffusion Models Hallucinate?
Text-to-image (T2I) diffusion models synthesize images by progressively denoising latent representations conditioned on text prompts. Throughout this iterative process, the model must simultaneously satisfy three core alignment objectives: conveying prompt semantics (SC), maintaining spatial plausibility (SA), and ensuring factual correctness (KG). Each of these objectives imposes directional alignment pressure on the generative trajectory, collectively forming the Hallucination Tri-Space $\mathcal{T}^{3}$.
To model alignment dynamics within this space, we define a real-time cognitive tension vector:
@@
which captures the evolving alignment demands along each cognitive axis at time step $t$ for prompt $p$. When the generative trajectory fails to regulate these tensions, hallucinations emerge—not as random outliers, but as structured projection shifts resulting from excessive load or directional misbalance in cognitive constraints.
We identify two key indicators of tension-induced risk:
@@
To quantify how alignment tension imbalance translates into hallucinated outputs, we formalize the generative deviation as a trajectory drift $\Delta\vec{t}$—a directional shift in the trajectory caused by cumulative cognitive misregulation. Specifically, as the model progresses through denoising steps, unresolved tensions along SC, SA, and KG axes distort the latent dynamics, pushing the trajectory away from the prompt-aligned manifold $\mathcal{M}_{\text{ideal}}$.
To make this projection shift interpretable and tractable, we model it as a trajectory drift component induced by cumulative anisotropic tension. The additional force $\Gamma(\vec{\tau})\cdot n(z)$ represents a directional deviation from the ideal trajectory, where $\Gamma(\vec{\tau})$ is a learned mapping from the alignment risk vector $\vec{\tau}=[\tau_{sc},\tau_{sa},\tau_{kg}]$ to a perturbation coefficient that scales with both the magnitude and imbalance of tension. A higher magnitude $|\vec{\tau}|$ indicates elevated alignment stress, while larger variance among $\tau_{i}$ components reveals tension anisotropy.
In this view, hallucinations are not caused by abrupt noise but by a persistent, directionally biased shift in the generative trajectory, where one or more tension components dominate and steer the diffusion process into biased latent regions outside the prompt-aligned manifold $\mathcal{M}_{\text{ideal}}$. We formalize this drift as:
@@ 5 Alignment Risk Code: Modeling Hallucination Tensions in T2I Diffusion
To quantitatively model the internal forces that drive projection shifts during image generation, we introduce the Alignment Risk Code (ARC), a dynamic tension vector that encodes the instantaneous alignment pressures along key cognitive axes. This section first formalizes the generative space as a tri-orthogonal cognitive field, then defines how ARC captures both the magnitude and directionality of tension-induced deviations.
@@
5.1 Hallucination Tri-Space: Semantic, Structural, and Knowledge Tensions
@@
Here, $\mathcal{A}_{i}(z_{t},p)$ is a scalar potential function reflecting how well latent state $z_{t}$ aligns with the $i$th cognitive goal. A large $\tau_{i}$ implies strong restorative pressure in subspace $\mathcal{M}_{i}$.
Figure 3 visualizes this tension space. Ideal generation maintains low and balanced $\tau_{i}$ values, preserving trajectory alignment with $\mathcal{M}_{\text{ideal}}$. However, when $\vec{\tau}$ becomes both large in norm and skewed in distribution, the trajectory bends toward a biased submanifold, leading to hallucinations aligned with dominant tension directions.

### v3

@@ Abstract
Despite remarkable progress in image quality and prompt fidelity, text-to-image (T2I) diffusion models continue to exhibit persistent “hallucinations”, where generated content subtly or significantly diverges from the intended prompt semantics. While previous works often regarded them as unpredictable artifacts, we argue that these failures reflect deeper, structured misalignments within the model’s generation process. In this work, we reinterpret hallucinations as trajectory drift within a latent alignment space. By tracking internal representations over time and analyzing diffusion trajectories across diverse prompts, we discover that hallucinated samples consistently deviate along structured paths that cluster into three separable failure modes. These emergent clusters correspond to distinct cognitive tensions: semantic coherence, structural alignment, and knowledge grounding. We then formalize this three-axis space as the Hallucination Tri-Space and introduce the Alignment Risk Code (ARC): a dynamic vector representation that quantifies real-time alignment tension during generation. The magnitude of ARC captures overall misalignment, its direction identifies the dominant failure axis, and its imbalance reflects tension asymmetry. Based on this formulation, we develop the TensionModulator (TM-ARC): a lightweight controller that operates entirely in latent space. TM-ARC monitors ARC signals and applies targeted, axis-specific interventions during the sampling process. Extensive experiments on standard T2I benchmarks demonstrate that our approach significantly reduces hallucination without compromising image quality or diversity. This framework offers a unified and interpretable approach for understanding and mitigating generative failures in T2I systems.
@@ 1 Introduction
Figure 1: Visualizing hallucination as trajectory drift in latent alignment space in T2I models. Successful generations (middle row) follow a coherent sampling trajectory from noise to data, remaining close to the prompt intent. Hallucinatory generations (bottom row) exhibit tension drift, leading to semantic and structural deviations.
@@
We begin by revisiting a foundational question in the generative process of diffusion models: Are hallucinations merely stochastic artifacts of sampling, or do they reflect deeper structural failures rooted in internal tension dynamics? To eliminate ambiguity and isolate the structural source of hallucinations, we analyze the latent generation trajectories of diverse T2I samples. See section 1.1, the finding reveals that hallucinations are not random outliers, but emerge along structured directions in the generative manifold and T2I generation is not a static mapping but a dynamic traversal through a latent tension space. In this process, each prompt implicitly defines a multiaxial cognitive tension field, and the diffusion process must iteratively traverse this space while balancing competing alignment forces. We term this dynamic imbalance cognitive alignment tension, which unfolds along the denoising trajectory. When one or more tension axes dominate, they disrupt the generative equilibrium, leading to a trajectory drift $\Delta\vec{t}$ that manifests as hallucination. Ideally, the process should stay within the balanced alignment manifold $\mathcal{M}_{\text{ideal}}$, but the imbalance in alignment tensions disrupt this path, pulling the model away from intended semantics.
@@
This realization leads us to a formal abstraction despite their surface diversity, hallucinations often stem from imbalances along three core alignment axes. We define the Hallucination Tri-Space $\mathcal{T}^{3}$, a latent space structured by orthogonal tensions:(1) Semantic Coherence (SC): the alignment between prompt entities and generated object categories; (2) Structural Alignment (SA): the fidelity of spatial layout and positional relationships; (3) Knowledge Grounding (KG): the factual and commonsense plausibility of generated content. Rather than classifying hallucinations, $\mathcal{T}^{3}$ models tension dynamics explicitly. By tracking the shift $\Delta\vec{t}$, it enables quantitative monitoring of imbalance during generation, providing a principled basis for real-time diagnosis and control. To quantify the cognitive alignment tension, we project the trajectory drift $\Delta\vec{t}$ onto the three axes of the Hallucination Tri-Space, yielding a real-time vector we term the Alignment Risk Code (ARC): $\vec{\tau}(p,t)=[\tau_{\text{SC}}(p,t),\tau_{\text{SA}}(p,t),\tau_{\text{KG}}(p,t)]^{\top}$. This ARC vector models the live state of alignment tensions during sampling. Its magnitude reflects the overall cognitive stress, its imbalance indicates the degree of tension asymmetry, and its direction characterizes the dominant source of misalignment. Rather than diagnosing hallucinations after generation, ARC enables in-process monitoring of when and how generative deviation begins to emerge. To leverage this signal for intervention, we introduce the TensionModulator (TM-ARC), a lightweight controller that operates fully within the latent space. TM-ARC interprets ARC dynamics and injects targeted, axis-specific corrections through three specialized submodules: (1) SC-Gate: mitigates semantic drift; (2) SA-Tuner: restores spatial integrity; (3) KG-Augment: reinforces factual grounding. Each submodule is activated with adaptive weights derived from the ARC vector, enabling fine-grained, real-time adjustment of the generative trajectory. This ensures the model remains within the ideal alignment manifold $\mathcal{M}_{\text{ideal}}$ throughout the denoising process, significantly reducing hallucinations without compromising image diversity or quality. The persistence of hallucinations stems not from prompt complexity but from the accumulation of misalignments across multiple cognitive axes. We model this phenomenon using the ARC vector, which captures rising tensions in dimensions such as knowledge grounding and semantic coherence. Leveraging this signal, our TM-ARC controller actively intervenes during generation, steering the trajectory back toward the ideal manifold $\mathcal{M}_{\text{ideal}}$. Our contributions are threefold:
@@
1.1 Empirical Motivation: A Constructive Probe Study
Before introducing the ARC Tri-Space framework, we conduct a probing experiment to investigate whether hallucinations in T2I diffusion models stem from random stochasticity or from structured failures rooted in latent misalignment. We hypothesize that hallucinations arise from a phenomenon we term trajectory drift, where the generation path diverges from the manifold of faithful samples due to accumulating alignment tensions. (See Appendix for implementation details).
Setup.
We sample $N=10$ high-fidelity, hallucination-free images per prompt from DrawBench Saharia et al. (2022), using them to construct a time-indexed success manifold $\mathcal{M}_{t}$ in latent space. We measure Mahalanobis distance between each hallucinated sample $x^{f}_{t}$ and $\mathcal{M}_{t}$ over time. Across 600 hallucinated cases, 78.3% exhibit statistically significant deviation ($Z$-score >> 3.0), with an average bifurcation point at $t_{b}=12.4\pm 3.7$.
Observations.
(1) Failure modes cluster naturally: Latent vectors at bifurcation points yield strong 3-cluster structure ($\text{ARI}=0.71$, $\text{NMI}=0.68$), with top-2 PCA components explaining 91.2% variance. (2) Cluster visualization: t-SNE projections (Figure 2) show well-separated groups corresponding to semantic, structural, and knowledge deviations. A Gaussian noise baseline yields no structure. (3) Consistency: These modes recur across 94.7% of prompts and show stable directional drifts. Bifurcation points vary by type: $t_{b}^{SC}=8.2$, $t_{b}^{SA}=14.8$, $t_{b}^{KG}=18.3$.
From Trajectories to Tensions.
These findings confirm that hallucinations are not random artifacts, but structured deviations along tension axes in latent space. Motivated by this, we define a real-time alignment vector $\bm{\tau}=[\tau_{SC},\tau_{SA},\tau_{KG}]$ to quantify projection drift along the ARC Tri-Space. This vector forms the basis for our design in ARC and TM-ARC, enabling not just detection but proactive intervention.
@@ 2 Why Do T2I Diffusion Models Hallucinate?
Text-to-image (T2I) diffusion models synthesize images by progressively denoising latent representations conditioned on text prompts. Throughout this iterative process, the model must simultaneously satisfy three core alignment objectives: conveying prompt semantics (SC), maintaining spatial plausibility (SA), and ensuring factual correctness (KG). Each of these objectives imposes directional alignment pressure on the generative trajectory, collectively forming the Hallucination Tri-Space $\mathcal{T}^{3}$. To model alignment dynamics within this space, we define a real-time cognitive tension vector:
@@
which captures the evolving alignment demands along each cognitive axis at time step $t$ for prompt $p$. When the generative trajectory fails to regulate these tensions, hallucinations emerge—not as random outliers, but as structured projection shifts resulting from excessive load or directional misbalance in cognitive constraints. We identify two key indicators of tension-induced risk:
@@
To quantify how alignment tension imbalance translates into hallucinated outputs, we formalize the generative deviation as a trajectory drift $\Delta\vec{t}$—a directional shift in the trajectory caused by cumulative cognitive misregulation. Specifically, as the model progresses through denoising steps, unresolved tensions along SC, SA, and KG axes distort the latent dynamics, pushing the trajectory away from the prompt-aligned manifold $\mathcal{M}_{\text{ideal}}$. To make this projection shift interpretable and tractable, we model it as a trajectory drift component induced by cumulative anisotropic tension. The additional force $\Gamma(\vec{\tau})\cdot n(z)$ represents a directional deviation from the ideal trajectory, where $\Gamma(\vec{\tau})$ is a learned mapping from the alignment risk vector $\vec{\tau}=[\tau_{sc},\tau_{sa},\tau_{kg}]$ to a perturbation coefficient that scales with both the magnitude and imbalance of tension. A higher magnitude $|\vec{\tau}|$ indicates elevated alignment stress, while larger variance among $\tau_{i}$ components reveals tension anisotropy. In this view, hallucinations are not caused by abrupt noise but by a persistent, directionally biased shift in the generative trajectory, where one or more tension components dominate and steer the diffusion process into biased latent regions outside the prompt-aligned manifold $\mathcal{M}_{\text{ideal}}$. We formalize this drift as:
@@ 3 Alignment Risk Code: Modeling Hallucination Tensions in T2I Diffusion
To quantitatively model the internal forces that drive projection shifts during image generation, we introduce the Alignment Risk Code (ARC), a dynamic tension vector that encodes the instantaneous alignment pressures along key cognitive axes. This section first formalizes the generative space as a tri-orthogonal cognitive field, then defines how ARC captures both the magnitude
@@
and directionality of tension-induced deviations.
@@
Here, $\mathcal{A}_{i}(z_{t},p)$ is a scalar potential function reflecting how well latent state $z_{t}$ aligns with the $i$th cognitive goal. A large $\tau_{i}$ implies strong restorative pressure in subspace $\mathcal{M}_{i}$. Figure 3 visualizes this tension space. Ideal generation maintains low and balanced $\tau_{i}$ values, preserving trajectory alignment with $\mathcal{M}_{\text{ideal}}$. However, when $\vec{\tau}$ becomes both large in norm and skewed in distribution, the trajectory bends toward a biased submanifold, leading to hallucinations aligned with dominant tension directions.